OpenStudy (anonymous):

what's the property of 8+3/5=3/5+8 1=1/3*3 4*0=0 N(ca)= (nc)a

OpenStudy (anonymous):

commutative property of addition a+b=b+a zero property of multiplication anything multiplied by zero is equal to zero associative property of multiplication a(bc)= (ab)c

OpenStudy (mathstudent55):

Each equation represents one property. The zero property of multiplication is a * 0 = 0 The commutative property of addition is a + b = b + a The multiplicative inverse property is a * 1/a = a, for a =/= 0 The associative property of multiplication is (ab)c = a(bc)